OpenStudy (anonymous):

Explain why (4,1) is not a solution to the equation y=3x+1

OpenStudy (anonymous):

when x=4 in the equation why does not =1

OpenStudy (anonymous):

because the 4 in the coordinate is x and 4*3 =1 +13 not 1

OpenStudy (anonymous):

3(4)+1=13